OpenStudy (anonymous):

Find the inverse of y=log base2 (x-5)+2.

Nnesha (nnesha):

first step swap x and y \[\huge\rm x =\log_2 (y-5)+2\] now try to solve for y

Nnesha (nnesha):

move the 2 to the left side and then convert log to exponential form

OpenStudy (anonymous):

I'm getting stuck after 2^log base2 (y-5) which is equal to y-5

Nnesha (nnesha):

i'm sorry what do you mean by 2^log ?

OpenStudy (anonymous):

so yes technically it looks like this x+2 = y-5 then x = y-3.

OpenStudy (anonymous):

2 to the power of log_2 (y-5) is equal to y - 5 after raising 2 to each side

Nnesha (nnesha):

well you have to move 2 to the left side so subtract 2 both sides \[\huge\rm x =\log_2 (y-5)+2\] \[\huge\rm x-2=\log_2(y-5)\] right ? now convert log to exponential form

Nnesha (nnesha):

the*

Nnesha (nnesha):

 that's how you should convert log to exponential form